# LSPRAG: LSP-Guided RAG for Language-Agnostic Real-Time Unit Test Generation

### v1

@@ 1. Introduction
Unit testing for modern software systems is crucial yet labor-intensive (Zhao et al., 2017; Beller et al., 2015). To ensure reliability, test suites must achieve high coverage to expose subtle bugs and edge cases. It requires developers to have a deep understanding of the internal logic and dependencies of the codebase, including how the functions under test interact with their surrounding implementation. This challenge is even greater in today’s software development, where enterprise codebases frequently span multiple programming languages, and developers expect real-time test generation as they code (Abidi et al., 2019; Yang et al., 2024; Li et al., 2021; Mayer et al., 2017).
Recent advances in Large Language Models (LLMs) have invigorated research on real-time multi-language unit-test generation. Commercial systems such as GitHub Copilot (Thakkar, 2022) already offer one-click test generation, boosting test quality with simple heuristics—for example, giving higher weight to files a developer has recently opened (Friedman, 2021). At the same time, Repository-level Retrieval-Augmented Generation (RAG) techniques supply LLMs with additional code context obtained via textual similarity (Lu et al., 2022; Zhang et al., 2023), graph-based relationships (Liu et al., 2024b, d), or even web search (Zhang et al., 2024b; Xia et al., 2024).
@@
Figure 1. Unit test generation scenario for focal method checkout. To cover the “true” block of the first branch of the focal method, the generator needs the definition of isValid from another file.:
Existing RAG approaches for code generation, unfortunately, also struggle to discover such a use-definition relationship precisely. This is because the RAG’s embedding-and-similarity approach depends on superficial cues such as similar function names, variable names, and comments to infer relationships. These cues are often noisy and prone to variation across coding styles and conventions. Recent research (Liu et al., 2024d; Cheng et al., 2024) recognizes this deficiency and augments retrieval with static-analysis-based searches; however, that solution demands substantial human effort and remains tied to specific programming languages, limiting their generalizability in today’s software development. This highlights the urgent need for a new approach for real-time and language-agnostic unit test generation.
@@
(2) Hard to guarantee valid tests in real-time. Even when concise and relevant context is collected, simply providing LLMs with more context does not guarantee the generation of valid test code. As noted in prior research (Zhang et al., 2024a; Liu et al., 2024c; Zhang et al., 2024c), large code context slices often introduce noise, which can obscure the original intent and reduce the LLM’s ability to generate correct code. As a result, the generated test code is likely to contain syntax errors, which hinder high-coverage test generation rather than support it. Previous works (Xie et al., 2023; Lemieux et al., 2023; Fraser and Arcuri, 2011; Pacheco and Ernst, 2007; Altmayer Pizzorno and Berger, 2024) have mitigated this issue by executing the generated test code in advance and collecting error messages in a fixing loop. However, this approach is not practical for a real-time scenario, as it requires a significant amount of time to compile and execute the generated test code. Furthermore, the codebase under test is often not compilable or executable in real-time.
@@ 2.1. Language Server Protocol
The Language Server Protocol (LSP) was introduced by Microsoft (Microsoft, 2024) to address a significant challenge in software development: the need for each editor and IDE to have a unique implementation for every programming language’s analysis features. As the number of editors and languages proliferated, this approach became unsustainable, forcing language authors to repeatedly rewrite the same analysis logic. At the same time, tool vendors struggled to keep pace with the emergence of new languages. LSP has resolved these problems by standardizing communication between a language server and any compliant editor or IDE.
@@ 5.1. Key Token Extraction
To overcome the limitations of the LSP, LspRag leverages the Abstract Syntax Tree (AST). The AST provides a language-agnostic structural view that is essential for understanding program hierarchy and can be easily implemented in a language-agnostic way, which is crucial for multi-language support. The effectiveness of this technique has been demonstrated in prior RAG research (Cheng et al., 2024; Saxena, 2023). LspRag reuses widely-used Tree-sitter (Max Brunsfeld and others, 2024) to construct a AST for focal method. The critical step in this process is to bridge the structural AST view with the lexical LSP tokens. To achieve this, LspRag preserves the precise source code location (i.e., line and column numbers) of every node during AST construction. This allows for a precise mapping between an AST node and its corresponding set of tokens if the node’s location span contains the tokens’ locations.
@@ 6. Implementation
To enhance usability, LspRag was implemented as a VS Code extension, which integrates seamlessly into the developer’s existing workflow and reduces the barriers to adoption. The tool leverages the VS Code Extension API, which provides comprehensive capabilities for user interface interactions, editor functionality, and command handling. To achieve language generalizability, we designed LspRag to be language-agnostic. It leverages the LSP to consume semantic information from any compliant language server and utilizes Tree-sitter (Max Brunsfeld and others, 2024) for robust source code parsing into an AST. Consequently, LspRag is compatible with any programming language supported by an LSP server and Tree-sitter. We have validated LspRag’s functionality with several major language servers, including Pylance for Python (Microsoft, 2024), the Oracle Extension Pack for Java (Oracle, 2024), and the official Go extension (Go Team at Google, 2024).
@@ 7.1. Experiment Setup
Benchmarks. Our evaluation is conducted on six open-source projects written in Python, Java, and Golang. Python and Java were selected due to their common use in prior unit test generation research. Golang was chosen as it is widely adopted in industry but less examined in academic studies, which we anticipated would present different challenges for LLMs. As shown in Table 2, we selected two projects per language, all of which are frequently used as benchmarks in the unit test generation domain (Xie et al., 2023; Wang et al., 2024b, 2023; Lemieux et al., 2023; Lukasczyk and Fraser, 2022).
@@
  • CodeQA (Saxena, 2023): An open-source RAG pipeline designed for code generation. Its approach, which involves indexing the codebase and parsing code structure with an AST, is analogous to that used in modern AI-assisted IDEs like Cursor (Cursor, 2025).
@@ 8. Discussion
Threats to Validity. We identify two primary external threats to the generalizability of our approach. First, the effectiveness of LspRag is inherently coupled with the correctness of the underlying LSP server implementation for a given language. Our approach relies on the LSP server to provide accurate static analysis results (e.g., definitions, references, and type hierarchies). Consequently, if a programming language lacks robust LSP support, or for servers with erratic or incomplete analysis capabilities, the quality of the retreived context would be degraded, potentially impacting the quality of the generated code. Second, the behavior of LspRag can be influenced by environment-specific configurations. For example, the behavior of LSP server is affected by different IDEs. We observed that for a Java project, VS Code (Microsoft, 2025) might define the class path as the project’s top-level directory, whereas Cursor (Cursor, 2025) sets it to the src/ subdirectory. Such discrepancies can affect which files are analyzed, potentially affecting the LSP’s static analysis results. Our prototype accommodates these variations through configuration.
@@ 9. Related Work
Existing Approaches on Unit Test Generation. Automated unit test generation has been a long-standing goal in software engineering, with two primary waves of innovation: classical search-based and symbolic techniques, and more recent LLM-based approaches. Classical techniques primarily focus on maximizing a specific criterion, most often code coverage. Search-Based Software Testing (SBST) tools (Deljouyi et al., 2025; Fraser and Arcuri, 2011; Lin et al., 2020; Pacheco and Ernst, 2007; Sapozhnikov et al., 2024) for Java, and tools (Lukasczyk and Fraser, 2022; Lemieux et al., 2023; Altmayer Pizzorno and Berger, 2024) for Python, employ evolutionary algorithms or random testing to generate test suites. While powerful, these methods are inappropriate for real-time. They rely on whole-program compilation, heavy instrumentation, and numerous execution iterations—steps that are too slow and fragile when the codebase is in flux, changing every few minutes during active development. Their operation assumes a stable, buildable project state, which is often not the case in the early, exploratory stages of writing new code. More recently, LLMs have been applied to test generation, often outperforming classical tools in terms of the readability and initial quality of generated tests (Xie et al., 2023; Dakhel et al., 2024; Wang et al., 2024a; Altmayer Pizzorno and Berger, 2024; Ryan et al., 2024). However, most of these works have selected an iterative loop either for error detection or coverage feedback. For instance, Altmayer et al. (Altmayer Pizzorno and Berger, 2024) propose coverage-guided pipelines where an LLM generates tests, which are then executed to gather coverage data; this feedback is used to prompt the LLM to cover the remaining branches. While these approaches have improved previous work in terms of latency and coverage, they are still inappropriate for real-time scenarios. Furthermore, both classical and many LLM-based approaches are tailored for specific programming languages (e.g., Java and Python), limiting their generalizability.
@@
Retrieval-Augmented Generation. RAG (Lewis et al., 2020b; Karpukhin et al., 2020) mitigates LLM’s limitation of having finite training data. Initially proposed for NLP, it has been widely applied to code, notably through repository-level techniques. These use code-specific features for context retrieval, such as AST-based chunking (Saxena, 2023; Ding et al., 2023; Shou et al., 2023) or program-analysis-based retrieval (Zhang et al., 2024d; Liu et al., 2024b, d; Chen et al., 2022). Repository-level RAG techniques are representative of this effort. They used code-specific features to retrieve context, such as AST-based chunking (Saxena, 2023; Ding et al., 2023; Shou et al., 2023), or program-analysis based retrieval (Zhang et al., 2024d; Liu et al., 2024b, d; Chen et al., 2022). These techniques expose a fundamental trade-off. AST-based chunking is multi-lingual, but its reliance on superficial features leads to imprecise retrieval. In contrast, program-analysis-based retrieval achieves far greater precision by understanding the code’s execution and data-flow relationships. However, this precision comes at a steep price: they require building and maintaining complex, language-specific static analyzers, thereby limiting their applicability to one specific language (Cheng et al., 2024; Liu et al., 2024d, b). We compare LspRag with DraCo(Cheng et al., 2024) and CodeQA(Saxena, 2023) as they are the only open-sourced, referred code-aware RAG works.
@@
Neuro-Symbolic Approaches in Software Engineering. Recent neuro-symbolic approaches combine LLMs with program analysis for enhanced code intelligence. This synergy aids code auditing and bug detection, with tools like LLMDFA (Wang et al., 2024c) and DeepConstr (Go et al., 2024) using LLMs for dataflow and constraint generation while symbolic techniques validate paths and mitigate hallucinations. For unit test generation, Tratto (Daka et al., 2025) uses a neural module to propose tokens and a symbolic module to constrain the search space based on program context. However, a common limitation is that they are tailored to a single programming language. In contrast, LspRag addresses this gap by leveraging the LSP to perform language-agnostic neuro-symbolic unit test generation.

### v2

@@ 1. Introduction
Unit testing for modern software systems is crucial yet labor-intensive (Zhao et al., 2017; Beller et al., 2015). To ensure reliability, test suites must achieve high coverage to expose subtle bugs and edge cases. It requires developers to have a deep understanding of the internal logic and dependencies of the codebase, including how the functions under test interact with their surrounding implementation. This challenge is even greater in today’s software development, where enterprise codebases frequently span multiple programming languages, and developers expect real-time test generation as they code (Abidi et al., 2019; Yang et al., 2024b; Li et al., 2021; Yang et al., 2024a).
Recent advances in Large Language Models (LLMs) have invigorated research on real-time multi-language unit-test generation. Commercial systems such as GitHub Copilot (Thakkar, 2022) already offer one-click test generation, boosting test quality with simple heuristics—for example, giving higher weight to files a developer has recently opened (Friedman, 2021). At the same time, Repository-level Retrieval-Augmented Generation (RAG) techniques supply LLMs with additional code context obtained via textual similarity (Lu et al., 2022; Zhang et al., 2023a), graph-based relationships (Liu et al., 2024d, c), or even web search (Zhang et al., 2024b; Xia et al., 2024).
@@
Figure 1. Unit test generation scenario for focal method checkout. To cover the “true” block of the first branch of the focal method, the generator needs the definition of isValid from another file.
Unit test generation scenario for focal method checkout. To cover the “true” block of the first branch of the focal method, the generator needs the definition of isValid from another file.
Existing RAG approaches for code generation, unfortunately, also struggle to discover such a use-definition relationship precisely. This is because the RAG’s embedding-and-similarity approach depends on superficial cues such as similar function names, variable names, and comments to infer relationships. These cues are often noisy and prone to variation across coding styles and conventions. Recent research (Liu et al., 2024c; Cheng et al., 2024) recognizes this deficiency and augments retrieval with static-analysis-based searches; however, that solution demands substantial human effort and remains tied to specific programming languages, limiting their generalizability in today’s software development. This highlights the urgent need for a new approach for real-time and language-agnostic unit test generation.
@@
(2) Hard to guarantee valid tests in real-time. Even when concise and relevant context is collected, simply providing LLMs with more context does not guarantee the generation of valid test code. As noted in prior research (Zhang et al., 2024a; Liu et al., 2024b; Zhang et al., 2024c), large code context slices often introduce noise, which can obscure the original intent and reduce the LLM’s ability to generate correct code. As a result, the generated test code is likely to contain syntax errors, which hinder high-coverage test generation rather than support it. Previous works (Xie et al., 2023; Lemieux et al., 2023; Fraser and Arcuri, 2011; Pacheco and Ernst, 2007; Altmayer Pizzorno and Berger, 2024) have mitigated this issue by executing the generated test code in advance and collecting error messages in a fixing loop. However, this approach is not practical for a real-time scenario, as it requires a significant amount of time to compile and execute the generated test code. Furthermore, the codebase under test is often not compilable or executable in real-time.
@@ 2.1. Language Server Protocol
The Language Server Protocol (LSP) was introduced by Microsoft (Microsoft, 2024a) to address a significant challenge in software development: the need for each editor and IDE to have a unique implementation for every programming language’s analysis features. As the number of editors and languages proliferated, this approach became unsustainable, forcing language authors to repeatedly rewrite the same analysis logic. At the same time, tool vendors struggled to keep pace with the emergence of new languages. LSP has resolved these problems by standardizing communication between a language server and any compliant editor or IDE.
@@ 5.1. Key Token Extraction
To overcome the limitations of the LSP, LspRag leverages the Abstract Syntax Tree (AST). The AST provides a language-agnostic structural view that is essential for understanding program hierarchy and can be easily implemented in a language-agnostic way, which is crucial for multi-language support. The effectiveness of this technique has been demonstrated in prior RAG research (Cheng et al., 2024; Saxena, 2023). LspRag reuses widely-used Tree-sitter (tree-sitter, 2024) to construct a AST for focal method. The critical step in this process is to bridge the structural AST view with the lexical LSP tokens. To achieve this, LspRag preserves the precise source code location (i.e., line and column numbers) of every node during AST construction. This allows for a precise mapping between an AST node and its corresponding set of tokens if the node’s location span contains the tokens’ locations.
@@ 6. Implementation
To enhance usability, LspRag was implemented as a VS Code extension, which integrates seamlessly into the developer’s existing workflow and reduces the barriers to adoption. The tool leverages the VS Code Extension API, which provides comprehensive capabilities for user interface interactions, editor functionality, and command handling. To achieve language generalizability, we designed LspRag to be language-agnostic. It leverages the LSP to consume semantic information from any compliant language server and utilizes Tree-sitter (tree-sitter, 2024) for robust source code parsing into an AST. Consequently, LspRag is compatible with any programming language supported by an LSP server and Tree-sitter. We have validated LspRag’s functionality with several major language servers, including Pylance for Python (Microsoft, 2024b), the Oracle Extension Pack for Java (Oracle, 2024), and the official Go extension (Go Team at Google, 2024).
@@ 7.1. Experiment Setup
Benchmarks. Our evaluation is conducted on six open-source projects written in Python, Java, and Golang. Python and Java were selected due to their common use in prior unit test generation research. Golang was chosen as it is widely adopted in industry but less examined in academic studies, which we anticipated would present different challenges for LLMs. As shown in Table 2, we selected two projects per language, all of which are frequently used as benchmarks in the unit test generation domain (Xie et al., 2023; Wang et al., 2024c, 2023; Lemieux et al., 2023; Lukasczyk and Fraser, 2022).
@@
  • CodeQA (Saxena, 2023): An open-source RAG pipeline designed for code generation. Its approach, which involves indexing the codebase and parsing code structure with an AST, is analogous to that used in modern AI-assisted IDEs like Cursor (Anysphere, 2025).
@@ 8. Discussion
Threats to Validity. We identify two primary external threats to the generalizability of our approach. First, the effectiveness of LspRag is inherently coupled with the correctness of the underlying LSP server implementation for a given language. Our approach relies on the LSP server to provide accurate static analysis results (e.g., definitions, references, and type hierarchies). Consequently, if a programming language lacks robust LSP support, or for servers with erratic or incomplete analysis capabilities, the quality of the retreived context would be degraded, potentially impacting the quality of the generated code. Second, the behavior of LspRag can be influenced by environment-specific configurations. For example, the behavior of LSP server is affected by different IDEs. We observed that for a Java project, VS Code (Microsoft, 2025) might define the class path as the project’s top-level directory, whereas Cursor (Anysphere, 2025) sets it to the src/ subdirectory. Such discrepancies can affect which files are analyzed, potentially affecting the LSP’s static analysis results. Our prototype accommodates these variations through configuration.
@@ 9. Related Work
Existing Approaches on Unit Test Generation. Automated unit test generation has been a long-standing goal in software engineering, with two primary waves of innovation: classical search-based and symbolic techniques, and more recent LLM-based approaches. Classical techniques primarily focus on maximizing a specific criterion, most often code coverage. Search-Based Software Testing (SBST) tools (Deljouyi et al., 2025; Fraser and Arcuri, 2011; Lin et al., 2020; Pacheco and Ernst, 2007; Sapozhnikov et al., 2024) for Java, and tools (Lukasczyk and Fraser, 2022; Lemieux et al., 2023; Altmayer Pizzorno and Berger, 2024) for Python, employ evolutionary algorithms or random testing to generate test suites. While powerful, these methods are inappropriate for real-time. They rely on whole-program compilation, heavy instrumentation, and numerous execution iterations—steps that are too slow and fragile when the codebase is in flux, changing every few minutes during active development. Their operation assumes a stable, buildable project state, which is often not the case in the early, exploratory stages of writing new code. More recently, LLMs have been applied to test generation, often outperforming classical tools in terms of the readability and initial quality of generated tests (Xie et al., 2023; Dakhel et al., 2024; Wang et al., 2024b; Altmayer Pizzorno and Berger, 2024; Ryan et al., 2024). However, most of these works have selected an iterative loop either for error detection or coverage feedback. For instance, Altmayer et al. (Altmayer Pizzorno and Berger, 2024) propose coverage-guided pipelines where an LLM generates tests, which are then executed to gather coverage data; this feedback is used to prompt the LLM to cover the remaining branches. While these approaches have improved previous work in terms of latency and coverage, they are still inappropriate for real-time scenarios. Furthermore, both classical and many LLM-based approaches are tailored for specific programming languages (e.g., Java and Python), limiting their generalizability.
@@
Retrieval-Augmented Generation. RAG (Lewis et al., 2020b; Karpukhin et al., 2020) mitigates LLM’s limitation of having finite training data. Initially proposed for NLP, it has been widely applied to code, notably through repository-level techniques. These use code-specific features for context retrieval, such as AST-based chunking (Saxena, 2023; Li et al., 2025; Lu et al., 2021) or program-analysis-based retrieval (Zhang et al., 2023b; Liu et al., 2024d, c; Chen et al., 2022). Repository-level RAG techniques are representative of this effort. They used code-specific features to retrieve context, such as AST-based chunking (Saxena, 2023; Li et al., 2025; Lu et al., 2021), or program-analysis based retrieval (Zhang et al., 2023b; Liu et al., 2024d, c; Chen et al., 2022). These techniques expose a fundamental trade-off. AST-based chunking is multi-lingual, but its reliance on superficial features leads to imprecise retrieval. In contrast, program-analysis-based retrieval achieves far greater precision by understanding the code’s execution and data-flow relationships. However, this precision comes at a steep price: they require building and maintaining complex, language-specific static analyzers, thereby limiting their applicability to one specific language (Cheng et al., 2024; Liu et al., 2024c, d). We compare LspRag with DraCo(Cheng et al., 2024) and CodeQA(Saxena, 2023) as they are the only open-sourced, referred code-aware RAG works.
@@
Neuro-Symbolic Approaches in Software Engineering. Recent neuro-symbolic approaches combine LLMs with program analysis for enhanced code intelligence. This synergy aids code auditing and bug detection, with tools like LLMDFA (Wang et al., 2024a) and DeepConstr (Go et al., 2024) using LLMs for dataflow and constraint generation while symbolic techniques validate paths and mitigate hallucinations. For unit test generation, Tratto (Daka et al., 2025) uses a neural module to propose tokens and a symbolic module to constrain the search space based on program context. However, a common limitation is that they are tailored to a single programming language. In contrast, LspRag addresses this gap by leveraging the LSP to perform language-agnostic neuro-symbolic unit test generation.